two concentric spring arranged one within other, control the operation of valve,the particular of the outer spring are 12coils, 25mm coil diameter, wire diameter 3mm and intial compression when valve closed 5mm the intial compression of the inner spring 11mm and total load is 150N calculate the stiffness of each spring and wire diameter ofvinner spring is 2mm and mean diameter is 15mm, take G as 81.5Gpa
1
Expert's answer
2021-10-06T00:46:46-0400


A force of 2 kN is acting on a circular rod with diameter 25 mm. The stress in the rod can be

"σ = (20 103 N) / (π ((20 10-3 m) / 2)2)
 = 127388535\times 2 (N/m2) 
 = 127\times 2 (MPa)
= 381(MPa)"


The change of length can be calculated by:


"dl = \u03c3 l_o \/ E"

"= (381 106 Pa) (2 m) \/ (83 109 Pa) \n = 0.00381 m\n = 3.81 mm"